OpenStudy (anonymous):

Find the derivative of f(x)=x^3-2 using f'(x)= lim h->0 f(x+h)-f(x)/h.

OpenStudy (anonymous):

\[f(x)=x^3-2\\ f'(x)=\lim_{h\to0}\frac{\bigg((x+h)^3-2\bigg)-\bigg(x^3-2\bigg)}{h}\] Expand the numerator. You'll find that all the terms not containing \(h\) will disappear, leaving you with a few terms that do contain it.

OpenStudy (anonymous):

So take h out? and go from there?

OpenStudy (anonymous):

Not until you've expanded the numerator: \[(x+h)^3-2-(x^3-2)=x^3+3x^2h+3xh^2+h^3-2-x^3+2\]which simplifies to \[3x^2h+3xh^2+h^3\]

OpenStudy (anonymous):

where do you get the 3 from?

OpenStudy (anonymous):

It's a byproduct of the expansion: \[(x+h)^3=x^3+3x^2h+3xh^2+h^3\]

OpenStudy (anonymous):

okay so i take the h out and get \[3x ^{2}+3xh+h ^{2}\] where do i go from there?

OpenStudy (anonymous):

Yep, next thing is to take the limit: \[\lim_{h\to0}(3x^2+3xh+h^2)=\cdots\]

OpenStudy (anonymous):

so \[3x ^{2}+3x\]?

OpenStudy (anonymous):

Close, the second term also disappears, since it contains a factor of \(h\). You're left with just \(3x^2\).

OpenStudy (anonymous):

so it goes away even though there is an x there?

OpenStudy (anonymous):

Yep, the limit only takes \(h\) into consideration (since it is the limit as \(h\to0\), not \(x\)).

OpenStudy (anonymous):

You can treat \(x\) as any old constant in this context.

OpenStudy (anonymous):

Oh alright cool thank you!

OpenStudy (anonymous):

You're welcome